The package procIAGssymp.sty for the formatting of a paper with the style of the proceeding of symposia sponsored by IAG

B.Benciolini e-mail: battista.benciolini@ing.uitn.it

No Second Author, No Third Author No-Where Institute

Abstract. The package procIAGssymp.sty has been designed to help the formatting of a paper with the style of the proceeding of symposia sponsored by the ”International Association of Geodesy (IAG)” published by ”Springer-Verlag Berlin Heidelberg NewYork”. This paper itself is an example of the use of the package. It contains instructions and suggestions for the use of the package, while most of the comment included in procIAGssymp.sty are of a more technical nature.

1 Introduction

The package procIAGssymp.sty provides some (re-)definitions of L^{A}TEXcommands. The use of the package is explained in this paper. I have tried to provide a package that helps in the preparation of a paper that corresponds to the instructions distributed by the publisher of the proceedings of IAG-sponsored-symposia (i.e. by Springer-Verlag), but … please read the IMPORTANT NOTES inside the file

procIAGssymp.sty.

The most curious readers are invited to look directly at the source of this file and, more important, to the code in the file procIAGssymp.sty.

2 Suggestions for the user

2.1 General page lay-out and page numbering

The general lay-out of the pages is controlled by some assignment made by procIAGssymp.sty AND by the use of the class article that must be loaded with the proper options, i.e.

[a4paper,twocolumn,10pt] .

The package itself does not modify the page numbering with respect to the L^{A}TEXstandard. If it is necessary to avoid page numbering this can be obtained putting \pagestyle{empty} somewhere in the preamble of the file and

                                            \thispagestyle{empty} immediately after

\maketitle. If it is necessary to start page numbering from n, this can be obtained with

                                            \setcounter{page}{n}.

2.2 Top-matter

The title and the author(s) must be prepared

                                            with \title{ . . .} and \author{ . . .}

and they will be printed by \maketitle. The use of \thanks{…} inside the argument of \title{ …} is not appropriate because this command has been redefined. If there are several authors, their names must be separated by \and. The authors that share the same address must be listed consecutively and must be followed by \thanks{…} (with the address as argument). To see an example you can see the code of this file.

2.3 Sectioning

                                            The use of \section{. . . } and
                                            \subsection{. . . }, can be done as in

any L^{A}TEXpaper. These commands have been redefined so that they produce the first level and second level headers in the proper format. All the other sectioning commands have been redefined with a coherent style. The command

                                            \section*{References} can be used to start

the reference list. One of the following sections is just a test of the various sectioning commands.

3 The choice of the font

The standard Computer Modern fonts are used. I have not tried any other font.

4 Test of sectioning (1st level)

4.1 Test of sectioning (2nd level)

4.1.1 Test of sectioning (3rd level)

This is just an example of the use and

abuse of \section, \subsection and
\subsubsection. We can also test \paragraph
and \subparagraph.

Paragraph This is a paragraph, headed by

\paragraph.

Subparagraph This is a sub paragraph. And the section end HERE.

5 Technicalities

5.1 About the format of the top-matter

The main structure of the topmatter is governed by \@maketitle which is properly redefined. The command \thanks has been redefined to print the authors’ address immediately after the authors’ names. The new form of this command makes it unsuitable to put a note to the title.

5.2 Compatibility

The package has been used together with the article class and with several packages in-

cluding: makeidx, amsmath, amssymb, apalike,
array, epsfig, graphicx and verbatim. I have

not found any compatibility problem.

6 Recommendation

I recommend all the user of the package procIAGssymp.sty to send me any comment about it.

APPENDIX Other TEX and L^{A}TEX material

This is an appendix not directly related to the main topic of this pages. Some of the TEX and L^{A}TEX tools that I have prepared for personal use can be useful for others TEX-friends.

A plain-TeX output routine for multicolumn page can be found in my personal web

site http://ing.unitn.it/~bencioli. The file

is quite simple and it is internally documented.

A L^{A}TEX2e package that allows the re-use of the content of \title and \author anywhere in the document is available on CTAN at : . . .

/macro/latex/contrib/suppoted/rectopma

A L^{A}TEX2e class for multi-author volumes. The class collective.cls is a tool designed for the preparation of a document composed by several contributions written independently by different authors and binded together in the proper format by an editor. The typical applications are the preparation of the proceedings of a meeting or of an issue of a journal. The editor can do the job preparing a file, the main file that uses the document class collective and that loads the various papers, the contributions, that are prepared by the authors as standard (La)TeX files. A preliminary release of collective.cls is available on request. I will send it by e-mail upon request. It works with some limitations, but I think it is already a useful tool. The ”standard ”distribution requires doc-strip for installation; the unpacked set, including the class file, documentation and an example, is also available. (I have recently seen an other package with a similar pourpose posted in the CTAN. )